Brief Exercise 9-5 Corales Company acquires a delivery truck at a cost of $49,600. The truck is expected to have a salvage value of $3,800 at the end of its 5-year useful life. Assuming the declining-balance depreciation rate is double the straight-line rate.

Compute annual depreciation for the first and second years under the declining-balance method.
Answer #1

Double declining rate = 200/Useful life = 200/5 = 40%

Depreciation for first year

= 49,600 * 40%

= 19,840

Depreciation for second year

= (49,600 - 19,840) * 40%

= 11,904